OpenStudy (anonymous):

factor using the formula for the sum or difference of two cubes 8x^3-1 Please show steps.

OpenStudy (mathteacher1729):

Hint: Note that 8 = 2*2*2 or \(\large 8=2^3\)

OpenStudy (anonymous):

Is this \[8x^3-1 or (8x)^3-1? \]

OpenStudy (anonymous):

8x^3-1

OpenStudy (mathstudent55):

A difference of two cubes factors like this: \(a^3 - b^3 = (a - b)(a^2 + ab + b^2) \)

OpenStudy (mathstudent55):

For you, a = 2x, and b = 1.

OpenStudy (anonymous):

Formula for factoring difference of two cubes is a3 - b3 = (a - b) (a2 + ab + b2) consider a to be 2x , hence a3 is 8x3

OpenStudy (anonymous):

Yeah, @mathstudent55 's got it.

OpenStudy (anonymous):

Where did you get those numbers from?

OpenStudy (anonymous):

(2x-1)(4x^2+2x+1)

OpenStudy (anonymous):

Refer to mathstudent55's formula from above. 1.) 8x^3 - 1 = (2x)^3 - (1)^3 2.) Since a = 2x and b = 1, (2x-1) * [(2a^2) + (2x)(1) + (1)^2 3.) Simplify (2x-1) * [(2a^2) + (2x)(1) + (1)^2 --> (2x-1) * (4x^2 + 2x +1)

OpenStudy (anonymous):